Solve: y2(x-y)dx+x2(x-y)dy+z(x2+y2)dz=0
1
Expert's answer
2019-03-26T12:15:45-0400
"{dx \over {y^2(x-y)}}={dy \over {x^2(x-y)}}=-{dx \over {z(x^2+y^2)}}"

We chooze


"P_1={1 \over z}, Q_1={1 \over z}, R_1=x-y"

"{{y^2(x-y)} \over z}+{{x^2(x-y)} \over z}-{{(x^2+y^2)(x-y)} \over z}=0"

Then function u1(x, y, z) is determined as


"u_1(x, y, z)=\int{{1 \over z}dx}+\int{{1 \over z}dy}+\int{(x-y)dx}"

"{x \over z}+{y \over z}+(x-y)z=C_1"

"{dx \over {y^2(x-y)}}={dy \over {x^2(x-y)}}"

"{dx \over {y^2}}={dy \over {x^2}}"

"y^2dy=x^2dx"

"y^3=x^3+C_2"

The equations


"{x \over z}+{y \over z}+(x-y)z=C_1, y^3=x^3+C_2"

constitute the integral curves of the given differential equations.